# Higher memory effects and the post-Newtonian calculation of their gravitational-wave signals

### v1

@@ 6.1 Terminology for higher memory effects and calculation methods
We now summarize the computations required to compute the PN expressions for the moments of the news and the corresponding shear. We use the PN expressions for the radiative modes given in Sec. 9.5 of [30]. We reproduce the relevant modes at the orders at which they are required in B, for $m\geq 0$ [the negative $m$ values can be obtained using the relationship that $h_{l(-m)}=(-1)^{m}\bar{h}_{lm}$]. These modes are written in terms of the mass parameters $M$, $\eta$ and $\delta m$; the PN parameter $x$; and the gravitational-wave phase $m\phi(x)$ for a given spherical-harmonic mode indexed by $(l,m)$. To compute the electric-parity part of the quadrupole or octupole fluxes and pseudo-fluxes, we can directly substitute the relevant expressions for the modes in B into the expressions (5.2.1) and (5.2.1). For all the magnetic fluxes and pseudo-fluxes, as well as the $l\geq 4$ modes, we must calculate expressions analogous to (5.2.1) and (5.2.1) that give the spherical-harmonic modes of the scalar fluxes in terms of the radiative modes. The calculations proceed analogously to those in Sec. 5.2, and the results are also lengthy expressions, such as those found in Eqs. (5.2.1) and (5.2.1). We, thus, do not give the expressions for general PN radiative modes; rather, we give only the leading-PN results for compact binaries below.
@@
Instead of giving a large number of algebraic expressions of the form $c_{n+1}x^{n+1}e^{-im\phi}$ for the leading $(l,m)$-mode of the non-oscillatory and oscillatory contributions from the zeroth, first and second fluxes (and pseudo-fluxes), we summarize the expressions in three tables, Tables 12 and 3.
@@ 6.2 Post-Newtonian orders and coefficients of higher memory effects
${\mathcal{F}_{1}}$ $(2,0)$ $(3,3)$ $-\frac{509}{990}\sqrt{\frac{\pi}{21}}\delta m\eta^{2}$
@@
${\mathcal{G}^{\mathrm{rad}\,^{*}}_{2}}$ $(3,0)$ $(2,1)$ $-\frac{8749i}{1701}\sqrt{\frac{\pi}{10}}\delta m\eta^{2}$
@@ 7.2 Discussion and future directions
Much of the emphasis of the calculations in this paper was on determining the lowest multipole moments (quadrupole and octupole) both to compare with existing PN calculations and to determine the size of the new flux contributions to the moments relative to the leading part of the gravitational-wave strain for compact binary sources. This latter direction is a necessary pre-requisite for assessing the detection prospects of these new contributions to the higher moments of the news. Such forecasts for the lower moments in ground-based detectors have already appeared in the literature. Because the flux contributions are small compared to the dominant quadrupole modes, this is typically done by “stacking” [53]. For studies using the binary-black-hole mergers discovered to date by LIGO and Virgo [56, 57], the stacking involves combining the evidence ratio of models that include these contributions versus those that do not. For forecasts of measurement prospects, it is also useful to consider stacking the the signal-to-noise ratio of the memory signals in quadrature [53, 54], which approximates the total evidence ratio (for all merger events) with and without the higher memories (see [52] for more details about when this equivalence holds).
@@
As future work, it is natural to investigate if similar studies could be done for the new higher memory effects arising from the second fluxes or pseudo-fluxes. The results of this paper suggest that the oscillatory pieces of the non-radiative pseudo-flux would be the most promising contribution given its relatively low PN order (2.5 PN), and the fact that it appears in the dominant $l=2$, $m=2$ waveform mode in non-spinning compact binaries. Based on the facts that it has the same PN order as the spin memory effect and that the amplitudes of the effects have the same mass dependence and numerical order of magnitude, this suggests that it might also be detectable in the population of black-hole mergers by the next generation of ground-based GW detectors. To obtain the most accurate forecasts in this context, it would be better to use the full inspiral-merger-ringdown signals from numerical-relativity simulations rather than the inspiral-only PN signals considered in this paper, however.
@@ Appendix B Radiative modes for compact binaries in quasi-circular orbits
However, to compute the entries in Table 3 given by Eqs. (LABEL:eqn:_COM_U20) and (LABEL:eqn:_Grad*_U30), we needed to go to higher PN orders, because the leading terms cancel. There are subleading terms that that are out of phase by $\pi/2$ from the leading terms that do not cancel, however. These terms are related to radiation reaction effects, and the ones that we require appear at 1.5 PN order relative to the “Newtonian” mode. We give these terms the superscript “RR” for radiation reaction, and we list them below:
@@ Appendix D Tensors on the 2-sphere
where $\mathrm{I}=\mathrm{E},\mathrm{B}$. Deriving the lowering relationship takes a few more steps. Writing $\mathscr{D}^{B}(T^{\mathrm{I}}_{lm})_{BA_{1}\cdots A_{s-1}}=2m^{(B}\bar{m}^{C)}\mathscr{D}_{B}(T^{\mathrm{I}}_{lm})_{CA_{1}\cdots A_{s-1}}$, then using Eq. (D), the product rule, Eq. (D.8) (and its complex conjugate) and Eq. (D.10), allows us to derive the following expressions for lowering tensor harmonics:
@@
We now derive expressions for the contraction of two tensor harmonics of equal rank in terms of a sum of scalar harmonics. Using Eq. (D), we obtain

### v2

@@ 6.1 Terminology for higher memory effects and calculation methods
We now summarize the computations required to compute the PN expressions for the moments of the news and the corresponding shear. We use the PN expressions for the radiative modes given in Sec. 9.5 of [30]. We reproduce the relevant modes at the orders at which they are required in B, for $m\geq 0$ [the negative $m$ values can be obtained using the relationship that $h_{l(-m)}=(-1)^{m}\bar{h}_{lm}$]. These modes are written in terms of the mass parameters $M$, $\eta$ and $\delta m$; the PN parameter $x$; and the gravitational-wave phase $m\phi(x)$ for a given spherical-harmonic mode indexed by $(l,m)$. To compute the electric-parity part of the quadrupole or octupole fluxes and pseudo-fluxes, we can directly substitute the relevant expressions for the modes in B into the expressions (5.15) and (5.16). For all the magnetic fluxes and pseudo-fluxes, as well as the $l\geq 4$ modes, we must calculate expressions analogous to (5.15) and (5.16) that give the spherical-harmonic modes of the scalar fluxes in terms of the radiative modes. The calculations proceed analogously to those in Sec. 5.2, and the results are also lengthy expressions, such as those found in Eqs. (5.15) and (5.16). We, thus, do not give the expressions for general PN radiative modes; rather, we give only the leading-PN results for compact binaries below.
@@
We next summarize how we will present the results of our calculations. The signals associated with the higher memory effects can all be written schematically as
$r\sqrt{2}h_{lm}^{\mathrm{(HME)}}=\hat{h}_{lm}x^{k+1}e^{-im\phi},$ (6.1)
where $\hat{h}_{lm}$ is a coefficient proportional to the radiative modes $U_{lm}$ or $V_{lm}$ associated with the higher memory effects, with the phase $e^{-im\phi}$ and the leading power of $x$ factored out. Because we compute the leading-order effect for the zeroth, first and second fluxes and pseudo-fluxes, for both electric and magnetic parities, for both oscillatory and non-oscillatory effects, and for both the moment of the news and the associated shear, there would be nearly 40 different expressions of the form in Eq. (6.1) to present. We do not write these expressions explicitly. Instead, we summarize the expressions in three tables. Table 1 contains the leading $l$ and $m=0$ non-oscillatory mode for each flux type, the corresponding PN orders [the power of $k$ in Eq. (6.1)], and the oscillatory modes that contribute to the higher memory effects; Table 2 contains the same information for the oscillatory modes. Finally, Table 3 focuses just on the shear associated with the higher memory effects, and has the coefficient $\hat{h}_{lm}$ for the leading flux or pseudo-flux. For the modes in Table 3, it is then straightforward to reconstruct the corresponding radiative moments $U_{lm}$ or $V_{lm}$ [or $h_{lm}$ via Eq. (5.6)] from the results in Tables 13 and Eq. (6.1). Similarly, the polarizations $h_{+}$ and $h_{\times}$ could be computed using the expression for $h_{lm}$ and Eq. (5.5) using the conventions for the spin-weighted spherical harmonics defined in D.
We also do not plot the higher-memory signals as a function of time or of the PN parameter $x$. At leading Newtonian order, the time dependence of $x(t)$ and of the phase $\phi(x)$ are given by
$x(t)=\frac{1}{4}\left[\frac{\eta}{5M}(t_{c}-t)\right]^{-1/4},\qquad\phi(x)=-\frac{x^{-5/2}}{32\eta}.$ (6.2)
Thus, the flux or pseudo-flux contributions will look no different from a higher PN correction to a given $(l,m)$ spherical harmonic mode of the waveform. Namely, the oscillatory signals will be chirp signals with a slowly increasing amplitude and frequency; the non-oscillatory signals will be just slowly increasing amplitudes. The more complex features of these memory signals will arise during the last stages of the inspiral, the merger and the ringdown; however, the PN approximation does not hold in this regime, so we are unable to plot the waveform from the later stages of the merger using the PN results in this paper. Instead, see [24], particularly Figure 3, for explicit plots of some of these contributions in a numerical relativity simulation, which covers the late inspiral, merger and ringdown.
@@ 6.2 Post-Newtonian orders and coefficients of higher memory effects
${\mathcal{F}_{1}}$ $(2,0)$  [^10] The expression for this coefficient is somewhat lengthier than the others. It is given by $\frac{2}{315M}\sqrt{\frac{\pi}{15}}\eta^{2}\biggr{[}80M^{2}(-121+538\eta)-3\delta m^{2}(23+20\ln 2)\biggr{]}.$ (6.3) The reason for the longer expression is that the leading-order terms in the modes given in Table 1 vanish, but subleading terms do not. These subleading terms all appear at 2.5 PN order relative to the quadrupole moment, and they are out of phase by $\pi/2$ with the leading modes (because they arise from dissipative, radiation-reaction effects). The product of these subleading terms with the leading moments does not cancel. $(3,3)$ $-\frac{509}{990}\sqrt{\frac{\pi}{21}}\delta m\eta^{2}$
@@
${\mathcal{G}^{\mathrm{rad}\,^{*}}_{2}}$ $(3,0)$  [^11] This expression is also longer, for the same reasons discussed in the previous footnote. It is given by $\frac{512}{25}\sqrt{\frac{\pi}{105}}(424\eta-97){M\eta}^{3}.$ (6.4) $(2,1)$ $-\frac{8749i}{1701}\sqrt{\frac{\pi}{10}}\delta m\eta^{2}$
@@ 7.2 Discussion and future directions
Much of the emphasis of the calculations in this paper was on determining the lowest multipole moments (quadrupole and octupole) both to compare with existing PN calculations and to determine the size of the new flux contributions to the moments relative to the leading part of the gravitational-wave strain for compact binary sources. These leading-order results could be generalized to include sub-leading corrections from higher PN terms in the oscillatory waveforms. The assumption of non-spinning compact objects could also be revisited to determine additional subleading corrections that enter from including spins. As one of the main goals was to compute the leading-order effects to simply determine which of the many flux terms merit further study, we will leave calculations with spinning PN results (which are more involved computationally) to future studies.
Understanding the largest higher-memory flux terms is a pre-requisite for assessing the detection prospects of these new contributions to the higher moments of the news. Such forecasts for the lower moments in ground-based detectors have already appeared in the literature. Because the flux contributions are small compared to the dominant quadrupole modes, this is typically done by “stacking” [53]. For studies using the binary-black-hole mergers discovered to date by LIGO and Virgo [56, 57], the stacking involves combining the evidence ratio of models that include these contributions versus those that do not. For forecasts of measurement prospects, it is also useful to consider stacking the signal-to-noise ratio of the memory signals in quadrature [53, 54], which approximates the total evidence ratio (for all merger events) with and without the higher memories (see [52] for more details about when this equivalence holds).
@@
As future work, it is natural to investigate if similar studies could be done for the new higher memory effects arising from the second fluxes or pseudo-fluxes. The results of this paper suggest that the oscillatory pieces of the non-radiative pseudo-flux would be the most promising contribution given its relatively low PN order (2.5 PN), and the fact that it appears in the dominant $l=2$, $m=2$ waveform mode in non-spinning compact binaries. Based on the facts that it has the same PN order as the spin memory effect and that the amplitudes of the effects have the same mass dependence and numerical order of magnitude, this suggests that it might also be detectable in the population of black-hole mergers by the next generation of ground-based GW detectors. For example, since the spin memory has a peak strain on the order of $10^{-24}$ for a GW150914-like event, and an SNR in Advanced LIGO of order 0.01 [14], it is not unreasonable to suppose the same would be true for this new radiative pseudo-flux term. To obtain the most accurate forecasts in this context, it would be better to use the full inspiral-merger-ringdown signals from numerical relativity simulations rather than the inspiral-only PN signals considered in this paper. This is an important reason why we leave this effort for future work.
@@ Appendix B Radiative modes for compact binaries in quasi-circular orbits
However, to compute the entries in Table 3 given by Eqs. (6.3) and (6.4), we needed to go to higher PN orders, because the leading terms cancel. There are subleading terms that that are out of phase by $\pi/2$ from the leading terms that do not cancel, however. These terms are related to radiation reaction effects, and the ones that we require appear at 1.5 PN order relative to the “Newtonian” mode. We give these terms the superscript “RR” for radiation reaction, and we list them below:
@@ Appendix D Tensors on the 2-sphere
where $\mathrm{I}=\mathrm{E},\mathrm{B}$. Deriving the lowering relationship takes a few more steps. Writing $\mathscr{D}^{B}(T^{\mathrm{I}}_{lm})_{BA_{1}\cdots A_{s-1}}=2m^{(B}\bar{m}^{C)}\mathscr{D}_{B}(T^{\mathrm{I}}_{lm})_{CA_{1}\cdots A_{s-1}}$, then using Eq. (D.6), the product rule, Eq. (D.8) (and its complex conjugate) and Eq. (D.10), allows us to derive the following expressions for lowering tensor harmonics:
@@
We now derive expressions for the contraction of two tensor harmonics of equal rank in terms of a sum of scalar harmonics. Using Eq. (D.6), we obtain